set up an equation and solve.
find three even consecutive integers such that twice the largest is equal to two more than three times the smallest.

Answer by rothauserc(4718)

let x be the first even integer, then
:
x, x+2, x+4 represent our three consecutive integers
:
we know that
:
2(x+4) = 3x + 2
:
2x+8 = 3x + 2
:
x = 6
:
*****************************************
Our three consecutive even integers are
6, 8, 10
:
check our answer
:
2(10) = 3(6) + 2
:
20 = 20
:
Our answer checks
*****************************************
: